# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

### v1

@@ VI Conclusion
In this paper, we considered the joint optimization of assigning SF, TP, BW, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G wireless link to optimize the global system EE in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.

### v2

@@ V-A Simulation setup
To satisfy the LoRa quality of service constraint, probabilistic path loss exponents for LoS and NLoS are given by $\delta_{\text{LoS}}=2$ and $\delta_{\text{NLoS}}=2.5$. We consider a Gaussian random variables $\mathcal{X}\sigma_{\text{LoS}}=5$ and $\mathcal{X}\sigma_{\text{NLoS}}=20$. In addtion, we consider a carrier frequency of $f=868\,\text{MHz}$. Furthermore, the agents can dynamically select configurations from three adjustable parameters. Particularly, SF is chosen from $\{7,8,9,10,11,12\}$, TP is selected from $\{2,5,8,11,14\}\,\text{dBm}$, and bandwidth $W$ configured from $\{125,250,500\}\,\text{kHz}$. For training, we employ a recurrent neural network (RNN) with 128 hidden units, a learning rate of $\alpha=3\times 10^{-4}$ for both actor and critic networks, and a soft target update coefficient of $0.01$. The training protocol uses a PPO clip range of $\epsilon=0.2$, a discount factor of $\gamma=0.99$, and the Adam optimizer.
@@
Figure 2: Centralized Critic and Decentralized Actor for MAPPO training.
@@ VI Conclusion
In this paper, we considered the joint optimization of assigning spreading factor, transmission power, bandwidth, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G propagation to optimize the global system EE in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.